OpenStudy (anonymous):

Write the partial fraction decomposition for a/x^(3)-x Please help!!

OpenStudy (anonymous):

oops is 1 instead of a

OpenStudy (primeralph):

What is the factorized form of (x^3)- x?

OpenStudy (anonymous):

\[\frac{ 1 }{ x^3-x }\] Taking x common from denominator \[\frac{ 1 }{x \left( x^2-1 \right) }\]